### v1

# Electrical conductivity of a random nanowire network: comparison of two-dimensional and quasi-three-dimensional models

Abstract
It is shown that the widely used two-dimensional model of random networks of metallic nanowires or carbon nanotubes significantly overestimates the number of contacts between elements compared to real systems, which, within the mean-field approach, leads to overestimated estimates of electrical conductivity, especially when the contact resistances between conductors make the main contribution to the electrical conductivity of the system. In the case of a two-dimensional model, the electrical conductivity of the system depends quadratically on the number density of conductors, whereas in the case of a three-dimensional model this dependence is linear.
Random networks of metallic nanowires (NWs) and carbon nanotubes (CNTs) on an insulating substrate attract researchers’ attention due to the use of such systems as components in various devices. A key issue is to identify the dependence of macroscopic physical properties of such networks (e.g., electrical conductivity) on the individual physical properties of their constituent elements (NWs, CNTs). While experimental studies reveal relationships between various macroscopic properties of the system as a whole, for example, the correlation between electrical conductivity and transparency of conducting films based on metallic NWs, it is very difficult, if at all possible, to experimentally obtain dependencies of macroscopic physical properties of the system on the properties of its constituent elements. To identify such dependencies, theoretical methods (e.g., dimensional analysis Benda, Cancès, and Lebental (2019)) and computer simulations Simoneau et al. (2013); Kim and Nam (2018) are used.
Since, for both NWs and CNTs, the typical aspect ratio (i.e., length-to-width ratio) is on the order of several hundred (see, e.g., Refs. Ustinovich et al., 2020; Wang et al., 2026), the model of one-dimensional conductors is widely used to simulate a random network of NWs or CNTs on an insulating substrate; i.e., it is assumed that conductors have zero width. The random network formed by such one-dimensional (1D) objects is considered two-dimensional (2D); a number of works are devoted to the study of the properties of such networks (see, e.g., Refs. Yi, Berhan, and Sastry, 2004; Heitz et al., 2011; Kim and Nam, 2018).
Let identical linear zero-width conductors of length be placed within a square domain of size () with periodic boundary conditions. The number density of conductors, i.e., the number of conductors per unit area, is
The centers of the conductors are assumed to be independently and identically distributed within , i.e., , where are the coordinates of the center of the conductor under consideration, while the orientations of the conductors follow a given orientational probability distribution function .
In a recent work Tarasevich, Eserkepov, and Vodolazskaya (2026), it was shown that, within the mean-field approach, the effective electrical conductivity of such a random network of conductors is
Here is the mean number of contacts per conductor,
where is the electrical resistance of a conductor, and is the resistance of each junction (intersection) between any conductors;
In the case of uniformly distributed orientations
and
In the case of large value of the number density of conductors and for
and for
(cf. Ref. Tarasevich, Eserkepov, and Vodolazskaya, 2022).
Although formula (2) was derived for a network in which all conductors lie strictly in one plane, it is easy to show that it remains valid even when the conductors form a layer of finite but small thickness; it is only necessary that the inclination of all conductors relative to a certain plane be small, which corresponds to real networks.
An alternative approach is to use a quasi-three-dimensional model (Q3D) Simoneau, Villeneuve, and Rochefort (2015); Forró et al. (2020); Daniels and Brown (2021). In this case, conductors are considered as thin but finite-diameter cylinders or spherocylinders. The properties of a network obtained by random deposition of such elongated objects of finite width onto a substrate are significantly different from the properties of a 2D network Daniels and Brown (2021). Figure 1 shows the distribution of the number of contacts per conductor for various values of the number density of conductors according to computer simulation data for rigid cylindrical conductors with length m and diameter nm Daniels and Brown (2021). Thus, the aspect ratio was 300.
Figure 2 shows the dependencies of the mean number of contacts per conductor on conductor number density for the 2D and Q3D models. While for a network of zero-width conductors a linear growth is observed, in the case of rigid finite-width conductors the dependence saturates.
Fundamentally different dependencies of the number of contacts per conductor on the number density of conductors in the 2D and Q3D models lead within a mean-field-approach to significantly different dependencies of electrical conductivity obtained on the number density (Fig. 3). In the case where contact resistances dominate over conductor resistance, the differences in electrical conductivities obtained within the 2D and Q3D models are most significant and for the values of the number density presented in the figure reach two orders of magnitude. These differences decrease significantly with increasing parameter .
It is clear that both models are only limiting cases of infinitely soft conductors (2D model) and infinitely rigid conductors (Q3D model). Real conductors possess some flexibility, so the number of contacts per conductor, network connectivity, and its electrical conductivity should occupy some intermediate position between these two limiting cases.
More realistic and, accordingly, more complex models were proposed in Ref. Simoneau, Villeneuve, and Rochefort, 2015; Colasanti et al., 2016. In Ref. Simoneau, Villeneuve, and Rochefort, 2015, CNTs were modeled as spherocylinders with soft shells; besides straight spherocylinders, objects consisting of several segments were used to reproduce the waviness of CNTs.
In Ref. Colasanti et al., 2016, networks of CNTs were simulated. A CNT was modeled as a cylinder that could bend upon contact. Each CNT was represented as a cylinder with radius and length chosen randomly according to some statistical distributions directly obtained from experimental measurements. The length of each CNT was determined from a log-normal distribution with a mean of 744 nm and a variance of 95 nm; the CNT diameter was constant and equal to 1.2 nm. Thus, the aspect ratio was 620. CNTs were placed sequentially one by one, with each time checking for intersections between the incoming CNT and the existing ones. The CNTs were placed on top, parallel to the substrate. For each individual CNT, the program used the obtained length and orientation relative to the plane to compute the coordinates of the two end points of the CNT itself. After obtaining the coordinates, the algorithm searches for all possible intersections with existing CNTs. A list of all such intersection points was generated, and among them, the point with the highest vertical coordinate was selected as the first intersection point. After this step, the original individual nanotube was split into two sub-segments, which were allowed to bend in the -direction. The algorithm again compiles a list of possible intersection points and selects the one with the largest coordinate. Two constraints were imposed on the algorithm, namely, a maximum bending angle (chosen to be less than ) and the CNT length, which remains constant throughout the bending process.
The authors of Ref. Colasanti et al., 2016 note that the 2D model overestimates the number of contacts between CNTs, even if the number of CNTs is very small. This artifact significantly impacts the reliability of the simulation, especially for dense films: since in the 2D model the network is more connected, there are more percolation paths; furthermore, since connections actually represent one of the main contributions to the film resistance, the 2D model cannot accurately reflect the real electrical behavior of the network.
In the case of a random three-dimensional (3D) packing of long rods, the average number of contacts per rod asymptotically in the limit of large length-to-width ratios () tends to Philipse (1996a, b); Williams and Philipse (2003). This result was confirmed in the paper Blouwolff and Fraden (2006), and in the case of the absence of compaction of the rod system the number of contacts is even less . The simulation results for the Q3D Daniels and Brown (2021) case are quite consistent with those obtained for the three-dimensional packing of long rods, but the average number of contacts per rod is even smaller. Thus, it can be assumed that when considering nanowires as rigid, elongated objects with a large length-to-thickness ratio, the number of contacts quickly saturates with increasing nanowire concentration and equals several (up to 10) contacts per nanowire.
When , the formula (2) becomes
and, in the case of equally probable conductor orientations, it reduces to
Thus, instead of the quadratic dependence of electrical conductivity on the nanowire concentration, which is given by the 2D approximation, we have a linear dependence.
It is clear that the real situation is much more complex; the dependence of the number of contacts per conductor on the number density of conductors significantly depends on our assumptions about the nature of interaction between conductor, the presence of a compressing (densifying) force, etc. If adhesion is excluded but gravity is considered significant, one can expect that the conductor will tend to occupy the lowest possible position, i.e., the supports will be located in hollows, leading to 4 contacts per newly added conductor. For example, the distributions presented in Ref. Daniels and Brown, 2021 for conductors with an aspect ratio of 300 indicate that the number of contacts can vary from 0 to 9 with a distribution maximum of approximately 3.
In any case, a network formed by straight thin conductors randomly placed on a substrate and having uniformly distributed orientations has significantly lower connectivity than a network formed by zero-width rods. However, in the case of a two-layer network formed by conductors oriented in two mutually perpendicular directions, the differences between the 2D and Q3D models are leveled out.
Acknowledgements.
The author thanks Avik Chatterjee for pointing out work Ref. Philipse, 1996a and Irina Vodolazskaya for careful reading of the manuscript and discussions.References
- Benda, Cancès, and Lebental (2019) R. Benda, E. Cancès, and B. Lebental, “Effective resistance of random percolating networks of stick nanowires: Functional dependence on elementary physical parameters,” Journal of Applied Physics 126, 044306 (2019).
- Simoneau et al. (2013) L.-P. Simoneau, J. Villeneuve, C. M. Aguirre, R. Martel, P. Desjardins, and A. Rochefort, “Influence of statistical distributions on the electrical properties of disordered and aligned carbon nanotube networks,” Journal of Applied Physics 114, 114312 (2013).
- Kim and Nam (2018) D. Kim and J. Nam, “Systematic analysis for electrical conductivity of network of conducting rods by Kirchhoff’s laws and block matrices,” Journal of Applied Physics 124, 215104 (2018).
- Ustinovich et al. (2020) K. B. Ustinovich, V. V. Ivanov, Y. M. Tokunov, A. A. Loshkarev, N. I. Sapronova, A. M. Vorobei, O. O. Parenago, and M. G. Kiselev, “Study of dispersions of carbon nanotubes modified by the method of rapid expansion of supercritical suspensions,” Molecules 25, 4061 (2020).
- Wang et al. (2026) J. Wang, J. Zuo, L. Ma, Y. Wang, J. Chen, Y. Li, and H. Zhang, “Rapid quantitative analysis of silver nanowire synthesis via UV-vis spectral deconvolution,” Journal of Applied Spectroscopy 92, 1387–1397 (2026), Russian Original: Zhurnal Prikladnoi Spektroskopii Vol. 92, No. 6, 832, 2025.
- Yi, Berhan, and Sastry (2004) Y. B. Yi, L. Berhan, and A. M. Sastry, “Statistical geometry of random fibrous networks, revisited: Waviness, dimensionality, and percolation,” Journal of Applied Physics 96, 1318–1327 (2004).
- Heitz et al. (2011) J. Heitz, Y. Leroy, L. Hébrard, and C. Lallement, “Theoretical characterization of the topology of connected carbon nanotubes in random networks,” Nanotechnology 22, 345703 (2011).
- Tarasevich, Eserkepov, and Vodolazskaya (2026) Y. Y. Tarasevich, A. V. Eserkepov, and I. V. Vodolazskaya, “Electrical conductivity of randomly placed linear wires: a mean field approach,” Nanotechnology 37, 045202 (2026).
- Tarasevich, Eserkepov, and Vodolazskaya (2022) Y. Y. Tarasevich, A. V. Eserkepov, and I. V. Vodolazskaya, “Electrical conductivity of nanorod-based transparent electrodes: Comparison of mean-field approaches,” Physical Review E 105, 044129 (2022).
- Simoneau, Villeneuve, and Rochefort (2015) L.-P. Simoneau, J. Villeneuve, and A. Rochefort, “Electron percolation in realistic models of carbon nanotube networks,” Journal of Applied Physics 118, 124309 (2015).
- Forró et al. (2020) C. Forró, S. J. Ihle, A. M. Reichmuth, H. Han, F. Stauffer, S. Weaver, A. Bonnin, M. Stampanoni, K. Tybrandt, and J. Vörös, “Visualizing and analyzing 3d metal nanowire networks for stretchable electronics,” Advanced Theory and Simulations 3, 2000038 (2020).
- Daniels and Brown (2021) R. K. Daniels and S. A. Brown, “Nanowire networks: how does small-world character evolve with dimensionality?” Nanoscale Horizons 6, 482–488 (2021).
- Colasanti et al. (2016) S. Colasanti, V. Deep Bhatt, A. Abdelhalim, and P. Lugli, “3-d percolative model-based multiscale simulation of randomly aligned networks of carbon nanotubes,” IEEE Transactions on Electron Devices 63, 1346–1351 (2016).
- Philipse (1996a) A. P. Philipse, “The random contact equation and its implications for (colloidal) rods in packings, suspensions, and anisotropic powders,” Langmuir 12, 1127–1133 (1996a).
- Philipse (1996b) A. P. Philipse, “The random contact equation and its implications for (colloidal) rods in packings, suspensions, and anisotropic powders,” Langmuir 12, 5971 (1996b).
- Williams and Philipse (2003) S. R. Williams and A. P. Philipse, “Random packings of spheres and spherocylinders simulated by mechanical contraction,” Physical Review E 67, 051301 (2003).
- Blouwolff and Fraden (2006) J. Blouwolff and S. Fraden, “The coordination number of granular cylinders,” Europhysics Letters (EPL) 76, 1095–1101 (2006).